# Generated Data with Fake Privacy: Hidden Dangers of Fine-tuning Large Language Models on Generated Data

### v1

@@ Abstract
Large language models (LLMs) have shown considerable success in a range of domain-specific tasks, especially after fine-tuning. However, fine-tuning with real-world data usually leads to privacy risks, particularly when the fine-tuning samples exist in the pre-training data. To avoid the shortcomings of real data, developers often employ methods to automatically generate synthetic data for fine-tuning, as data generated by traditional models are often far away from the real-world pertaining data. However, given the advanced capabilities of LLMs, the distinction between real data and LLM-generated data has become negligible, which may also lead to privacy risks like real data. In this paper, we present an empirical analysis of this underexplored issue by investigating a key question: "Does fine-tuning with LLM-generated data enhance privacy, or does it pose additional privacy risks?" Based on the structure of LLM’s generated data, our research focuses on two primary approaches to fine-tuning with generated data: supervised fine-tuning with unstructured generated data and self-instruct tuning. The number of successful Personal Information Identifier (PII) extractions for Pythia after fine-tuning our generated data raised over $20\%$. Furthermore, the ROC-AUC score of membership inference attacks for Pythia-6.9b after self-instruct methods also achieves more than $40\%$ improvements on ROC-AUC score than base models. The results indicate the potential privacy risks in LLMs when fine-tuning with the generated data.
@@ 1 Introduction
In recent times, large language models (LLMs) such as GPT-4 [1], LLaMA-3 [2], and Mistral [3] have demonstrated considerable success in text generation and have been extensively deployed for a variety of specific tasks, including chatbots and LLM agents. The impressive capabilities of these LLMs are largely attributed to the vast pre-training datasets sourced from the Internet or data providers, which may contain some sensitive data. Apart from that, the selected training data for LLMs will highly influence the model’s performance. Thus many LLM providers like OpenAI and Meta choose to keep their training data selection private. To assess the potential privacy risks on sensitive information or private training data, researchers have proposed numerous well-designed attacks associated with LLMs, such as membership inference attacks (MIA)[28, 36, 29, 11, 8], Personally Identifiable Information (PII) Attacks [23], and others [26, 29, 10]. Due to the huge amount of datasets used in LLMs’ training, these attacks are generally less effective compared to traditional models as studied in recent works [15, 11], resulting in a stronger sense of privacy for LLMs.
However, recent research [12] indicates that fine-tuning an LLM on datasets overlapping with the model’s pre-training data may pose privacy risks to the related parts of the pre-training data. In these studies, researchers fine-tune LLMs using a small subset of the pre-training data and found that this process also enhances the model’s memorization of other data points related to the fine-tuning data. The fine-tuning procedure can increase LLM’s memorization and elicit privacy risks, such as the potential extraction of sensitive information [23]. As a result, fine-tuning with real datasets carries potential privacy concerns, particularly when these datasets overlap with LLM’s pre-training data.
Like traditional machine learning, LLMs can also leverage generated data for fine-tuning. Notable examples include Alpaca [31] for instruction tuning and HH-RLHF [6] for preference optimization, among others. Moreover, researchers have developed various prompting techniques to help language models generate high-quality data for fine-tuning. For example, developers can provide short, human-written prefixes to LLMs, prompting them to complete the remaining content for fine-tuning. Another method involves prompting LLMs with input-output pairs related to specific seed tasks and guiding them to generate similar tasks with corresponding pairs for further fine-tuning. Like traditional machine learning, these generated datasets greatly improve the performance of language models and are widely adopted due to their flexibility and low data collection costs. This naturally raises the question: Does fine-tuning on entirely synthetic datasets generated by LLMs introduce privacy risks?
@@ 1.1 Threat Model
We investigate scenarios where LLM developers first train a model on their collected datasets and subsequently fine-tune specific LLMs for different domain-related tasks before publishing them online. Aware that fine-tuning these specialized LLMs using parts of the original training set could lead to private data leakage, developers opt to use generated data for this process. Our paper evaluates the potential risks that these fine-tuned models pose to the developers’ private data. Since most LLM developers only provide a query API for their models, potential attackers would be limited to querying the models to extract sensitive information. However, we also consider the more severe case where attackers can access the returned logits of the outputs. To assess privacy risks, we employ PII extraction and score-based MIA techniques.
@@ 1.2 Our Work
To explore the question, we begin by experimenting with fine-tuning language models using different types of generated data. We then employ MIA and PII attacks to evaluate potential privacy risks. Our study primarily examines the two most common fine-tuning scenarios for language models: supervised fine-tuning with unstructured data and instruction tuning, each targeting distinct goals.
@@
The first scenario is designed to enhance the model’s performance across various domains, such as improving comprehension or reasoning on specialized data. In this case, we prompt LLMs with task-specific prefixes and use their completions for fine-tuning. We denote the above procedure as supervised fine-tuning with unstructured data in our paper. In contrast, we follow the "self-instruct" methodology in the second scenario, which feeds input-output pairs related to predefined tasks into LLMs to generate similar tasks with new input-output pairs. Fine-tuning with this generated data not only enhances domain-specific capabilities but also improves the model’s ability to better follow user prompts. The details are listed as follows.
@@
Risks on Fine-Tuning with Unstructured Generated Data. Following the experiment’s setting in former work [12], we use the Enron email dataset to evaluate the potential privacy risks on fine-tuned Pythia models. We first use Pythia-12B to generate some email-related data, and then fine-tune the Pythia model with different model sizes on these generated datasets. After that, we do PII attacks following Wang et al.’s [33] setting on both the pre-trained model and fine-tuned models. The results demonstrate that supervised fine-tuning will amplify privacy risks even on the unstructured generated data (Section 3.3). After that, we ran experiments to analyze such privacy risks and found that the template and quality of generated data are the main factors that may influence PII’s success rate (Section 3.4).
@@
Risks on Fine-Tuning with Self-Instruct. In this section, we conduct experiments following the "self-instruct" tuning pipeline, as illustrated in previous research [34], to access potential privacy risks associated with Pythia’s pre-trained datasets. As the tasks related to the email datasets are limited and difficult to implement the self-instruct tuning, we opt for widely explored legal chatbot tasks for instruction fine-tuning for our task in this section. In line with self-instruct procedures, we initially designed several task descriptions and related input-output pairs (denoted as seed tasks) based on the FreeLaw datasets, the legally related subset of Pythia’s pre-training dataset. After that, we prompt LLama-3 with these seed tasks and ask it to procure the generated data for fine-tuning, including task descriptions, queries, and answers.
@@
Fine-tuning with the above data, we can get legal LLMs based on Pythia, which exhibit enhanced performance on legal question-answering tasks. Then we conduct the score-based MIA method following Duan et al. [15]’s setting, on the self-instruct tuned Pythia models and their pre-trained version. The results reveal that the AUC ROC score of MIA on FreeLaw datasets enjoys nearly $20\%$ improvement compared to the pre-trained model. These results indicate that the use of self-instructed data generated by large language models can slightly exacerbate the model’s privacy vulnerabilities. Further investigation reveals that the primary factor influencing the models’ privacy is the quality of the generated data.
We summarize our contributions as follows:
@@
  • Firstly, we evaluate the privacy risks of supervised fine-tuning in LLMs using generated data without an instructional structure, specifically through a PII attack. The results demonstrate that fine-tuning with generated email data increases the success rate of PII attacks by over $50\%$ compared to the pre-trained model. This suggests that training on generated raw data within the same domain can significantly amplify the privacy leakage associated with the LLM’s pre-training datasets.
@@
  • Secondly, we evaluate the privacy risks associated with LLMs fine-tuned using instruction-based data. Our analysis shows that self-instruct tuning on law-related tasks increases the model’s vulnerability within the Law dataset from its pre-training data. Specifically, the AUC-ROC score for a reference-based MIA attack on the fine-tuned Pythia-6.9b model rose by $20\%$ compared to the pre-trained model. These results suggest that self-instruct tuning can exacerbate privacy risks, especially in domains closely related to self-instruct tasks.
@@ 2 Preliminaries
In this section, we summarize LLM’s pre-training and various fine-tuning methods.
@@ 2.1 LLM’s Pre-training
Instead of training only on one domain of data during the training stage in traditional neural networks, LLMs usually take all possible texts from different domains, like various tasks (code and story) or multi-linguistic corpus. Furthermore, the pertaining data of LLM usually contains many unprocessed raw texts like HTML, and JSON files crawled directly from the Internet. Training with next-word prediction loss on these data of various domains and structures [35], LLMs build a general capability on text understanding and generation across different domains. Due to its capability in such
@@ 2.2 Fine-tuning Methods
Fine-Tuning with Unstructured Raw Texts. To extend LLMs’ abilities on various domain tasks, users can further fine-tune LLMs with different kinds of datasets, like some biomedicine and finance datasets. The fine-tuning procedure is similar with LLM’s pre-training but the data size will be smaller. After fine-tuning such specific datasets, LLMs’ ability in such domains will be enhanced and better for some professional tasks. Due to this reason, many researchers also propose methods to many close-sourced LLM providers like OpenAI also provide their API for fine-tuning with unstructured raw texts to help users train better
Instruct Tuning and Self-Instruct. Instruction tuning is a popular technique employed to enhance the ability of large language models (LLMs) to follow user prompts, thereby producing more accurate responses. In contrast to traditional fine-tuning, which utilizes raw textual data, instruction tuning necessitates the use of manually crafted instructions, user-generated prompts (inputs), and expected answers (outputs). Consequently, the collection of such data is heavily dependent on manual labeling, which is often resource-intensive. To address the challenges associated with gathering instruction data, researchers have proposed the "self-instruct" method. This approach involves using advanced LLMs to generate instruction-tuning samples, which can then be utilized for fine-tuning purposes.
To generate instruct data with good quality, self-Instruct first takes an initial dataset of instructions and their corresponding input-output examples, referred to as ’seed tasks’. For instance, an instruction may be ’What is the name of the victim in the following legal document?’ while input-output examples are some selected documents and mentioned victims in them. The quality and diversity of seed tasks are vital for the efficacy of the procedure. Once the seed tasks are ready, the rest of the procedure depends on the Generator, and the Target model, which are not necessarily distinct.
@@
  • Bootstrapping Examples. For each generated task, the generator generates new input-output examples by leveraging a similar few-shot prompting approach with instruction-input-output triples. Generated examples that are not in the desired form are excluded from the next step.
@@
LoRA and Quantization In 2021, Hu et al. proposed the Low Rank Adaptation (LoRA) method  [18] for efficient fine-tuning of larger models without utility loss. LoRA reduces occupied memory during fine-tuning by "freezing" a large portion of model parameters and updating the trainable parameters with low-rank approximation (i.e., adapter) of the update matrix. The adapter is optimized with respect to the loss function and multiplied by the scale factor before adding up to the target modules. This makes it possible to merge the base model with different adapters, that are much smaller in size compared to fully fine-tuned models.
To further reduce the computational cost of fine-tuning large models, Dettmers et al. introduced Quantized Low Rank Adaptation (QLoRA) method  [13]. QLoRA uses block-wise quantization for less precision loss and trains the quantized LoRA adapter.
@@ 2.3 Models
As the Pythia models’ training sets are open-sourced on the website, we choose this model as the target model to evaluate the potential privacy risks. We also use the powerful Llama-3 as the self-instruct method’s generator to generate high-quality fine-tuning data. Details for these models are listed as follows.
Pythia Suite Developed by EleutherAI, Pythia suite [7] provides open-source LLMs of sizes ranging from 14m to 12b. The models at each size are trained on both the standard and deduplicated version of The Pile [16].
In our work, Pythia models with sizes 410m, 1.4b, 2.8b, and 6.9b are used as target models while 12b model is used as the generator in section 3.
Llama-3-8b-Instruct is the smallest model in Meta’s open-source Llama-3 collection [5]. This model is chosen for its strong instruction-following capabilities and relatively compact size. It is utilized for creative generation tasks in section 4.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated significant success in various domain-specific tasks, with their performance often improving substantially after fine-tuning. However, fine-tuning with real-world data introduces privacy risks. To mitigate these risks, developers increasingly rely on synthetic data generation as an alternative to using real data, as data generated by traditional models is believed to be different from real-world data. However, with the advanced capabilities of LLMs, the distinction between real data and data generated by these models has become nearly indistinguishable. This convergence introduces similar privacy risks for generated data to those associated with real data. In this paper, we present an empirical analysis of this underexplored issue by investigating a key question: "Does fine-tuning with LLM-generated data enhance privacy, or does it pose additional privacy risks?" Our study investigates this question by examining the structural characteristics of data generated by LLMs, focusing on two primary fine-tuning approaches: supervised fine-tuning (SFT) with unstructured (plain-text) generated data and self-instruct tuning. In the scenario of SFT, the data is put into a particular instruction tuning format used by previous studies. We use Personal Information Identifier (PII) leakage and Membership Inference Attacks (MIAs) on the Pythia Model Suite and Open Pre-trained Transformer (OPT) to measure privacy risks. Notably, after fine-tuning with unstructured generated data, the rate of successful PII extractions for Pythia increased by over $20\%$, highlighting the potential privacy implications of such approaches. Furthermore, the ROC-AUC score of MIAs for Pythia-6.9b, the second biggest model of the suite, increases over $40\%$ after self-instruct tuning. Our results indicate the potential privacy risks associated with fine-tuning LLMs using generated data, underscoring the need for careful consideration of privacy safeguards in such approaches.
@@ 1 Introduction
Recently, large language models (LLMs) such as GPT-4 [1], LLaMA-3 [2], and Mistral [3] have demonstrated considerable success in text generation and have been extensively deployed for a variety of specific tasks, particularly as customized chatbots. The impressive capabilities of these LLMs are largely attributed to the vast pre-training datasets sourced from the Internet or data providers  [13]. The choice of training data plays a critical role in the performance of LLMs. As a result, many LLM providers, such as OpenAI and Meta, opt to keep their training data selection confidential. However, the training data often contains privacy-sensitive data from real individuals [12, 16]. To assess the potential privacy risks on sensitive information or private training data, researchers have proposed numerous well-designed attacks associated with LLMs, such as membership inference attacks (MIA)[45, 53, 46, 16, 14], Personally Identifiable Information (PII) attacks [38], and data extraction attacks [40, 46, 16].
However, recent research [17] highlights that fine-tuning an LLM on datasets overlapping with its pre-training data can pose privacy risks, especially to closely related pre-training portions. In [17], researchers fine-tuned LLMs using a small subset of the pre-training data and observed that this process also enhances the model’s memorization of other data points related to the fine-tuning data. The fine-tuning process can amplify an LLM’s memorization capabilities, potentially leading to privacy risks, such as the extraction of sensitive information [38]. Consequently, fine-tuning with real datasets raises significant privacy concerns, especially when these datasets overlap with the model’s pre-training data. However, no previous research has explored the implications of fine-tuning with generated data.
Similar to traditional machine learning, LLMs can also leverage generated data for fine-tuning. Notable examples include Alpaca [48] for instruction tuning and HH-RLHF [9] for preference optimization, among others. Moreover, researchers have developed various prompting techniques to help LLMs generate high-quality data for fine-tuning. For example, developers can use concise, human-written prompts to guide LLMs in generating content for fine-tuning. Alternatively, they can provide input-output pairs for specific seed tasks, which serve as prompts for the model to generate additional task-specific examples with corresponding pairs, facilitating further fine-tuning. These generated datasets greatly improve the performance of LLMs and are widely adopted due to their flexibility and low costs. This naturally raises the question: Does fine-tuning on entirely synthetic datasets generated by LLMs introduce privacy risks? This inquiry has not been addressed by previous research in terms of concrete privacy risks. Our findings suggest that despite common belief, generated data does not mitigate but exacerbates the risks.
@@ 1.1 Threat Model
We explore scenarios in which LLM developers initially train a model using their proprietary datasets and subsequently fine-tune it to perform various domain-specific tasks before making it publicly available. Recognizing the risk of private data leakage that may arise from fine-tuning specialized LLMs with portions of the original training set, developers opt to use generated data for this process. Our paper evaluates the potential privacy risks of fine-tuning using generated data. Since most LLM developers offer access to their fine-tuned models solely through a query-based API, potential attackers would be limited to querying the models to extract sensitive information. However, we also consider the more severe case in which attackers can access the returned logits of the outputs. To assess privacy risks, we employ PII extraction and score-based MIA techniques.
@@ 1.2 Our Work
To study the aforementioned risks, we begin by experimenting with fine-tuning LLMs using various types of generated data. We then employ MIA and PII attacks to evaluate the privacy risks. Our study primarily examines the two most common fine-tuning scenarios for language models: supervised fine-tuning (SFT) with unstructured data and instruction tuning.
The SFT with unstructured data scenario is designed to enhance the model’s performance across various domains, such as improving comprehension or reasoning on emails. In this case, we prompt LLMs with email-specific prefixes and use their completions for fine-tuning. This is explained in more detail in the Section 2.3.
@@
On the other hand, the self-instruct approach feeds existing tasks and input-output pairs into a capable LLM -such as GPT 3.5 as used by [52]- to generate similar tasks with new input-output pairs. Fine-tuning with this generated data not only enhances domain-specific capabilities but also improves the model’s ability to better follow user prompts. The details are listed as follows.
@@
Risks on Fine-Tuning with Unstructured Generated Data. Following the setting in former work [17], we use the Enron email dataset to evaluate the potential privacy risks on fine-tuned Pythia models. We first use Pythia-12b to generate an email dataset, and then fine-tune the Pythia model with different model sizes on these generated datasets. Then, we conduct PII attacks following Wang et al.’s [51] setting on both the pre-trained model and fine-tuned models. The results demonstrate that supervised fine-tuning amplifies privacy risks even in unstructured generated data (Section 3.3). After that, we ran experiments to analyze the privacy risks and found that the template and quality of generated data are the main factors that may influence PII’s success rate (Section 3.3.3).
In addition to these experiments, we conduct further evaluations using Facebook’s OPT model, which has no overlap with Enron or any email data in its training set. For this evaluation, we modify the Enron dataset to minimize overlap with the sensitive data in Pythia’s training data. These experiments not only extend the scope of our analysis but also serve to verify the results observed in our previous experiments, confirming the trends and insights derived from the Pythia-based evaluations in more challenging scenarios.
@@
Risks on Fine-Tuning with Self-Instruct. In Section 4 we conduct experiments following the "self-instruct" tuning pipeline, as illustrated in previous research [52]. Our aim is to examine the potential privacy risks associated with Pythia’s pre-trained datasets, The Pile [24]. We choose the FreeLaw [24] subset of The Pile for the privacy-sensitive nature of the law domain. In line with the self-instruct procedure, we initially designed 64 task descriptions focusing on legal expertise and 75 related input-output pairs (denoted as seed tasks) based on the FreeLaw dataset. After that, we prompt LLama-3 with these seed tasks to procure the generated data for fine-tuning, including task descriptions, related information, and answers.
@@
Fine-tuning Pythia models with the generated data, we can obtain LLMs that exhibit enhanced performance on legal question-answering tasks. Then we conduct the score-based MIA method following Duan et al. [21]’s setting, on the self-instruct tuned Pythia models and their pre-trained version. The results reveal that the AUC ROC score of MIA on FreeLaw datasets enjoys nearly $20\%$ improvement compared to the pre-trained model. These findings highlight that leveraging self-instructed data generated by LLMs can intensify the model’s susceptibility to privacy vulnerabilities. Further investigation reveals that the primary factor influencing the models’ privacy is the quality of the generated data. We summarize our contributions as follows:
@@
  • We evaluate the privacy risks of supervised fine-tuning in LLMs using generated data without an instructional structure, specifically through a PII attack. The results demonstrate that fine-tuning with generated email data increases the success rate of PII attacks by over $50\%$ compared to the pre-trained model. This suggests that training on generated raw data within the same domain can significantly amplify the privacy leakage associated with the LLM’s pre-training datasets.
@@
  • We evaluate the privacy risks associated with fine-tuned LLMs using instruction-based data. Our analysis shows that self-instruct tuning on law-related tasks increases the model’s vulnerability within the law-related subset from its pre-training data. Specifically, the AUC-ROC score for a reference-based MIA attack on the fine-tuned Pythia-6.9b model increased by $20\%$ compared to the pre-trained model. These results suggest that self-instruct tuning can exacerbate privacy risks, especially in domains closely related to self-instruct tasks.
@@ 2 Preliminaries
In this section, we summarize LLM’s prompting, pre-training, and various fine-tuning methods.
2.1 LLM Prompting
Prompting methods are strategies used to obtain specific responses from LLMs by designing the input text in particular ways. These methods can be used to increase the comprehension of LLMs on the task, often with purposes such as enhancing response quality or format. Some prompting methods depend on LLMs ability to recognize patterns and generate the response based on them. One such effective technique is few-shot prompting [13], where the prompt includes a handful of example input-output pairs that demonstrate the desired task. By presenting these examples, the LLM can infer the underlying task structure and produce appropriate responses to new inputs. For instance, to translate a particular text from English to German, the prompt might include "Hello" → "Hallo" and "Good morning" → "Guten Morgen," enabling the LLM to translate "Thank you" → "Danke."
@@ 2.2 LLM’s Pre-training
Pretraining large language models (LLMs) has significantly advanced with the development of transformer-based architectures, compared with former approaches like Word2Vec and GloVe. Notably, GPT-based models [42] pioneered the autoregressive pretraining paradigm, where the model learns to predict the next token in a sequence. GPT-2 [43] further demonstrated the capabilities of large-scale unsupervised learning, setting new benchmarks in various NLP tasks. Building on these successes, GPT-3 [13] introduced even larger models, with 175 billion parameters, and showcased remarkable performance across a wide range of tasks without requiring task-specific fine-tuning. More recently, open-sourced models such as LLaMA [50] have emerged, aiming to provide highly efficient alternatives by optimizing training and scaling strategies. LLaMA models, like GPT, are designed to excel in language understanding and generation while being more accessible for research and applications. Additionally, models like Pythia [11] and Mistral [3] have contributed to making large-scale autoregressive models available to a broader community, encouraging further exploration and refinement of pretraining techniques. Despite these advancements, challenges related to model bias, computational cost, and interpretability remain central to ongoing research in the field of LLM pretraining.
@@ 2.3 Fine-tuning Methods
Fine-Tuning with Unstructured Raw Texts. To enhance the capabilities of LLMs across various domain-specific tasks, users can fine-tune them further using specialized datasets, such as those from biomedicine, law, and finance domains. The fine-tuning process resembles the pre-training of LLMs but typically involves a smaller dataset. Because pre-trained LLMs already have a solid understanding of language intrinsics, fine-tuning them on domain-specific datasets can yield competitive performances. Additionally, several closed-source LLM providers, such as OpenAI, offer APIs for fine-tuning using unstructured raw text, enabling users to further optimize model performance for their specific needs.
Instruction Tuning and Self-Instruct. Instruction tuning is a popular technique employed to enhance the ability of large language models (LLMs) to follow user prompts, thereby producing more accurate responses. In contrast to traditional fine-tuning, which utilizes raw textual data, instruction tuning necessitates the use of manually crafted instructions, user-generated prompts (inputs), and expected answers (outputs). Consequently, the collection of such data is heavily dependent on manual labeling, which is often resource-intensive. To address the challenges associated with gathering instruction data, researchers have proposed the "self-instruct" method. This approach involves using advanced LLMs to generate instruction-tuning samples, which can then be utilized for fine-tuning purposes.
To generate instruct data with good quality, Self-Instruct first takes an initial dataset of instructions and their corresponding input-output examples, termed ’seed tasks’. For example, an instruction might be, "What is the name of the victim in the following legal document?" The corresponding input-output examples would consist of legal documents as inputs and the identified victims mentioned in them as outputs. The quality and diversity of seed tasks are vital for the efficacy of the procedure. Once the seed tasks are ready, the rest of the procedure depends on the Generator, and the Target model (see below), which are not necessarily distinct.
@@
  • Bootstrapping Examples. For each generated task, the generator creates new input-output examples using a similar few-shot prompting approach, incorporating instruction-input-output triples. Generated examples not in the desired form are excluded from the next step.
@@
Parameter-efficient fine-tuning (PEFT) methods. We describe below the PEFT methods used in this study:
LoRA and Quantization Hu et al. propose the Low Rank Adaptation (LoRA)  [27] for efficient fine-tuning of models without utility loss. LoRA reduces occupied memory during fine-tuning by "freezing" a large portion of model parameters and updating the trainable parameters with low-rank approximation (i.e., adapter) of the update matrix. The low-rank approximation involves decomposing a high-dimensional matrix into the product of two lower-dimensional matrices, reducing computational complexity. The update matrix refers to the changes applied to the original model parameters in each step. The adapter is optimized with respect to the loss function and multiplied by the scale factor to control the magnitude of the updates. This approach enables the integration of the base model with various adapters, which are significantly smaller in size compared to fully fine-tuned models. Besides the LoRA method, various methods are proposed to reduce LoRA’s parameter [32], increase safety [34] and etc [29].
To further reduce the computational cost of fine-tuning large models, Dettmers et al. introduced Quantized Low Rank Adaptation (QLoRA) method  [19]. QLoRA uses block-wise quantization which divides the model parameters into smaller blocks and quantizes each block separately. This technique minimizes precision loss and reduces computational overhead, enabling the training of the quantized LoRA adapter and we use QLoRA for Pythia experiments.
DoRA, an improvement over LoRA, was introduced in [37]. Unlike LoRA, which primarily focuses on low-rank adaptation, DoRA incorporates an additional trainable parameter: magnitude. This parameter enhances the model’s flexibility, enabling faster convergence while achieving higher precision during fine-tuning. We use DoRA to fine-tune the Facebook OPT 1.4b parameter model and the Pythia 2.8b parameter model on datasets that do not overlap with their training data.
@@ 2.4 Models
We choose Pythia and Open Pre-trained Transformer (OPT) language models as the target models to evaluate the potential privacy risks. We also use the powerful Llama-3 as the self-instruct method’s generator to generate high-quality fine-tuning data. Details for these models are listed as follows.
Pythia Suite is developed by EleutherAI [11] and provides open-source LLMs of varying sizes. The models at each size are trained on both the standard and deduplicated version of The Pile [24]. We use Pythia models with parameter sizes of 410m, 1.4b, 2.8b, and 6.9b as target models, while the 12b model serves as the generator in Section 3.
Llama-3-8b-Instruct is the smallest model in Meta’s open-source Llama-3 collection [5]. This model is chosen for its strong instruction-following capabilities and relatively compact size. It is used for creative generation tasks in Section 4.
OPT Language Models were introduced by Meta AI to provide researchers with access to high-performance language models [56]. These models are designed to approximately match the size and performance of the GPT-3 family of models. Pre-trained versions of OPT are available in various sizes, ranging from 125m to 66b parameters. In this work, we employ the 1.3 billion and 2.7 billion parameter versions of the OPT model. This choice was made because a subset of the Pile dataset  [24], including CommonCrawl, DM Mathematics, Project Gutenberg, HackerNews, OpenSubtitles, OpenWebText2, USPTO, and Wikipedia, was used in the training of these models. Notably, the Enron subset is excluded from this dataset. Consequently, the training data for the OPT models does not overlap with the data used in our method, ensuring data independence and mitigating potential biases.